OpenStudy (anonymous):

solve this system of equation..show steps please!! y^2=x^2-25 x^2-y^2=7

OpenStudy (anonymous):

No solution

OpenStudy (anonymous):

can u explain to me how u got tht?

OpenStudy (anonymous):

y^2 = x^2 - 25 y^2 - x^2 = -25 x^2 - y^2 = 7 multiply -1 on both sides y^2 - x^2 = -7 -7 does not equal to -25

OpenStudy (anonymous):

OHH...i was wondering why it wasnt working out!! thnk u!!!!

hero (hero):

y^2=x^2-25 y^2 = x^2-7 y^2 = y^2 x^2 - 25 = x^2 - 7 x^2 - x^2 = 25 - 7 0 =25 - 7 0 = 18 False

OpenStudy (anonymous):

thnk u too, hero

hero (hero):

:D